# Diffusion-Driven Generative Framework for Molecular Conformation Prediction

### 22 Dec 2023 v1

@@ Abstract
The task of inferring three-dimensional molecular configurations from their two-dimensional graph representations is of critical significance in the domains of computational chemistry and the development of pharmaceuticals. It contributes fundamentally to our grasp of molecular mechanisms and interactions. The rapid evolution of machine learning, especially in the realm of deep generative networks, has catalyzed breakthroughs in the precision of such predictive modeling. Traditional methodologies typically employ a bifurcated strategy: initially estimating interatomic distances followed by sculpting the spatial molecular structure via solving a distance geometry problem. This sequential approach, however, occasionally fails to capture the intricacies of local atomic arrangements accurately, thus compromising the integrity of the resultant structural models. Addressing these deficiencies, this work introduces an avant-garde generative framework: DDGF, which is predicated on the diffusion principles found in classical non-equilibrium thermodynamics. DDGF envisages atoms as discrete entities and is adept at guiding the reversal of diffusion morphing a distribution of stochastic noise back into coherent molecular forms through a process akin to a Markov chain. This transformation begins with the initial representation of a molecular graph in an abstract latent space, progressing to the realization of the three-dimensional forms via an elaborate bilevel optimization scheme, tailored to respect the task’s specific requirements. One of the formidable challenges in this modeling endeavor is the preservation of roto-translational invariance to ensure that the molecular conformations generated respect the laws of physics. Our proposed framework adeptly navigates this complexity, facilitating training from start to finish by fine-tuning a weighted variational lower bound concerning the conformation’s conditional probability. Extensive experiments validate the effectiveness of the proposed DDGF over the state-of-the-art methods.
@@ I Introduction
Graph representation learning has been highly successful in modeling molecules for various tasks, including property prediction [1, 2, 3, 4], time series prediction [5, 6] and molecule generation [7, 8]. In this approach, a molecule is typically represented as an atom-bond graph. Despite its effectiveness in various applications, a more intrinsic and informative representation for molecules is the 3D geometry, also known as conformation, where atoms are represented by their Cartesian coordinates. The 3D structures determine the biological and physical properties of molecules and hence play a key role in many applications such as computational drug and material design [9, 10, 11, 12]. Despite these advances, predicting stable molecular conformation remains a challenging problem. Traditional methods based on molecular dynamics (MD) or Markov chain Monte Carlo (MCMC) are computationally expensive and do not scale well to large molecules [13]. To overcome this challenge, researchers are exploring new methods that combine graph-based models with physics-based simulations or use reinforcement learning to optimize molecular conformations.
@@
Experimentally ascertaining the three-dimensional formations of molecules can be a daunting and resource-intensive process, characterized by both high costs and extended timeframes. Consequently, the quest for computationally deriving accurate and energetically favorable molecular structures has risen to prominence in the realm of computational chemistry. Traditional methodologies in this field have historically depended on techniques such as Markov chain Monte Carlo (MCMC) or molecular dynamics (MD) simulations [14], which involve proposing potential conformations and ascribing to them energy levels via either cost-effective empirical potential models or high-precision quantum chemical computations [15]. The tide has been shifting recently towards incorporating machine learning algorithms to estimate the conditional probability distribution $p(\mathbf{C}|G)$ for generating the stable conformations $\mathbf{C}$ of a given molecular graph $G$, building upon datasets of molecules whose stable conformations are already determined [16, 17]. Notably, two groundbreaking studies [17, 18] have introduced a sequential methodology. The initial phase involves predicting interatomic distances, which sets the stage for the subsequent creation of molecular structures in adherence to these predicted distances, a task that is facilitated by resolving a set of distance geometry challenges[19]. By harnessing the principles of distance geometry, these novel methods manage to capture the essential rotation and translation invariance characteristics of molecular structures, leading to their increasingly accurate predictions of molecular conformations.
@@
The sequential methodologies that first estimate interatomic distances and subsequently deduce conformations encounter a notable drawback: the risk of inaccuracies in local three-dimensional geometries. There are instances where the models for predicting distances might inadvertently lend credence to geometrically incompatible sets of distances, which can cascade into significant deviations. These inaccuracies, when input into the conformation generation stage, can yield atypical and implausible molecular structures. The root of such discrepancies can often be traced back to the distance prediction models, which prioritize the maximization of individual distance likelihoods rather than the holistic accuracy of the resulting conformations. In response to this challenge, our work introduces a streamlined, end-to-end framework that formulates conformation predictions directly from the molecular graph, eliminating the intermediate distance prediction phase. Our approach endeavors to holistically encapsulate the intricate dependencies and interplays among atoms, thus aspiring to elevate the precision and dependability of conformation prediction. This shift towards an integrated predictive model echoes the evolutionary trajectory seen in the domain of protein structure prediction. Innovations exemplified by the AlphaFold2 algorithm underscore the merit and impact of adopting such an end-to-end strategy. AlphaFold2’s refinement over its predecessor, AlphaFold, accentuates the benefits of consolidating distinct phases into a cohesive modeling process [20, 21]. Although AlphaFold2’s internal mechanisms remain partially undisclosed, its profound success in deducing protein structures from linear sequences underscores the potential of end-to-end methodologies in complex structural prediction tasks.
@@
In our research, we introduce DDGF, a cutting-edge, integrated framework tailored for the synthesis of molecular conformations via a sophisticated bilevel programming approach. Central to our method is the integration of roto-translational invariance within the conformational modeling scheme, wherein interatomic distances serve as pivotal variables. Our novel approach transcends the traditional distance prediction models that target minimization of prediction errors in the space of distances. By framing the problem within the context of bilevel optimization [22], DDGF concurrently addresses both the prediction of distances and the resolution of distance geometry challenges that are fundamental to conformational synthesis. To our knowledge, DDGF stands as the inaugural algorithm in the realm of molecular conformation synthesis that is trainable end-to-end while retaining the crucial aspects of roto-translational invariance. Our comprehensive experiments showcase DDGF’s exceptional capabilities, consistently outperforming the current leading methods across a suite of established benchmarks that span conformational synthesis and distance distribution analysis. Furthermore, our findings confirm the criticality of an end-to-end training objective in the creation of conformations that are not only realistic but also chemically significant.
@@ II Related Work
Conformation generation. Recent progress in the field of computational molecular modeling has spawned a plethora of innovative deep generative models and methodologies, each with its unique advantages and inherent challenges. For example, the CVGAE [16] leveraged a variational autoencoder (VAE) for the direct computation of 3D atomic structures. This model, however, did not adequately maintain the critical property of roto-translation equivariance, leading to suboptimal performance. To address this deficiency, newer models have turned to intermediate structural descriptors such as atomic distances and dihedral angles, known for their roto-translationally invariant nature—vital for the precise representation of molecular shapes[23]. Yet, this modular approach, which relies on these intermediate descriptors, introduces certain inefficiencies and complexities during the training phase and generating samples. Approaches like those proposed by [17], [18] involve predicting interatomic distance matrices and subsequently applying Distance Geometry solutions to derive spatial coordinates [19]. Although CONFVAE offered an advancement with its unified, bilevel optimization-based end-to-end framework [24], these techniques invariably encounter the problem of error magnification. The inaccuracies in distance estimations frequently lead to the misguidance of the coordinate determination mechanisms, culminating in the creation of molecular structures that are not just inaccurate but sometimes structurally unfeasible.
@@
To address this problem, CONFGF [25, 26] aimed to learn the gradient of log-likelihood in relation to coordinates. However, in practice, the model still relied on intermediate geometric elements. It estimated the gradient with respect to interatomic distances using denoising score matching (DSM) [27, 28] and then used the chain rule to derive the gradient of coordinates. Unfortunately, by learning the distance gradient via DSM, the model was trained with perturbed distance matrices that might violate the triangular inequality or contain negative values. Consequently, the model learned from invalid distance matrices but was tested with valid ones calculated from coordinates, leading to a serious out-of-distribution problem [29]. A recent work called GEOMOL [30] introduced a highly systematic rule-based pipeline. It focused on predicting a minimal set of geometric quantities, such as lengths and angles, and then reconstructed both local and global structures of conformations through a sophisticated procedure. Additionally, there have been efforts to utilize reinforcement learning for conformation search, exemplified by methods like TorsionNet [31]. However, this approach fundamentally differs from other existing methodologies as it relies on the rigid rotor approximation and only models torsion angles.
@@
Advances in Protein Topology Prediction. The landscape of protein topology prediction has expanded rapidly with a plethora of novel methodologies emerging. Innovations include the use of flow-based models [32], which adeptly map the conformations of protein main chains. Techniques employing recurrent neural architectures have been utilized for sequential amino acid modeling [33].  [34] introduced a novel approach using neural networks to emulate an energy landscape for inferring protein folds. The trailblazing AlphaFold algorithm [20, 35] has notably elevated the benchmark for protein structure prediction accuracy. It is crucial to note, however, that the structural simplicity of proteins, predominantly linear in nature, contrasts starkly with the complex, branched configurations of general molecules that include various ring structures. This structural disparity renders direct application of protein-folding methodologies less suitable for the intricate task of general molecular conformation prediction, underscoring the necessity for distinct modeling techniques that cater to the unique challenges presented by non-protein molecular structures.
@@
Point cloud generation. In the realm of three-dimensional structural synthesis, diffusion-driven models have made their mark, with pioneering efforts seen in recent literature [36, 37]. These approaches have principally targeted the generation of point clouds, a distinct domain from our inquiry. Point clouds typically lack the graph-like structure that embodies diverse atomic and bonding information, a cornerstone of our model’s domain. Furthermore, the principle of equivariance, although not extensively incorporated in these point cloud methodologies, is a cornerstone in our approach. This disparity in foundational concepts underscores the novelty of our model when juxtaposed against the landscape of existing methods.
@@
In summary, the landscape of deep generative models has ushered in a new era for the generation of molecular conformations. Each innovation offers unique benefits and faces distinct challenges. There is a continuous need for exploration to surmount the persisting hurdles in enhancing the precision and computational efficiency of these conformational generation techniques.
@@ III-D Bilevel Optimization
Bilevel programs are defined as optimization problems where a set of variables involved in the (outer) objective function are obtained by solving another (inner) optimization problem. Formally, given the outer objective function F and the inner objective H, and the corresponding outer and inner variables $\theta$ and $w$, a bilevel program can be formulated by:
@@
Bilevel programs have shown effectiveness in a variety of situations such as hyperparameter optimization, adversarial and multi-task learning, as well as meta-learning.
@@
where $\partial$ denotes the partial derivative to compute the Jacobian on immediate variables while $\nabla$ denotes a total derivative taking into account the recursive calls to $F$. The above gradient can be efficiently calculated by unrolling the optimization dynamics with back-propagation, i.e., reverse mode automatic differentiation  [43], where we repeatedly substitute $w_{\Phi,t}=\Phi(w_{\theta,t-1},\theta)$ and apply the chain rule.
@@ IV-A Diffusion Process
Drawing parallels from thermodynamics, we envision the molecules’ conformation C as particles within an ever-fluctuating system. As moments elapse, the pristine equilibrium state $\textbf{c}^{0}$ embarks on a stochastic journey through a series of increasingly disordered states $\textbf{c}^{t}$, culminating in a state of maximal entropy akin to white noise after $T$ transitional steps. This progression, distinctive from conventional latent variable constructs, is characterized in our diffusion model by a predetermined posterior distribution $q(\mathbf{c}^{1:T}|\mathbf{c}^{0})$, which remains constant throughout the learning process. Specifically, we define it as a Markov chain according to a fixed variance schedule $\beta_{1},\cdots,\beta_{T}$ :
@@
Note that, in this work we do not impose specific (invariance) requirement upon the diffusion process, as long as it can efficiently draw noisy samples for training the generative process $p_{\theta}(\mathbf{c}^{0})$.
Defining $\alpha_{t}=1-\beta_{t}$, and ${\bar{\alpha}}_{t}=\Pi^{t}_{s=1}\alpha_{s}$, we observe a noteworthy aspect of the forward process: the conditional distribution $q(\mathbf{c}^{t}|\mathbf{c}^{0})$ at any given $t$ can be precisely expressed as $q(\mathbf{c}^{t}|\mathbf{c}^{0})=\mathcal{N}(\mathbf{c}^{t};\sqrt{{\bar{\alpha}}_{t}}\mathbf{c}^{0},(1-{\bar{\alpha}}_{t})I)$. With an adequately extended series up to $T$ steps, this process inherently transforms $\mathbf{c}^{0}$ into a standard Gaussian distribution, implying a harmonious choice for $p(\mathbf{c}^{T}$ would be an isotropic Gaussian distribution.

### 21 Jan 2024 v2

@@ Abstract
The task of deducing three-dimensional molecular configurations from their two-dimensional graph representations holds paramount importance in the fields of computational chemistry and pharmaceutical development. This process significantly contributes to our comprehension of molecular mechanisms and interactions. The rapid advancement of machine learning, particularly within the domain of deep generative networks, has revolutionized the precision of predictive modeling in this context. Traditional approaches often adopt a two-step strategy: initially estimating interatomic distances and subsequently refining the spatial molecular structure by solving a distance geometry problem. However, this sequential approach occasionally falls short in accurately capturing the intricacies of local atomic arrangements, thereby compromising the fidelity of the resulting structural models. Addressing these limitations, this research introduces a cutting-edge generative framework named DDGF. This framework is grounded in the principles of diffusion observed in classical non-equilibrium thermodynamics. DDGF views atoms as discrete entities and excels in guiding the reversal of diffusion, transforming a distribution of stochastic noise back into coherent molecular structures through a process akin to a Markov chain. This transformation commences with the initial representation of a molecular graph in an abstract latent space, culminating in the realization of three-dimensional structures via a sophisticated bilevel optimization scheme meticulously tailored to meet the specific requirements of the task. One of the formidable challenges in this modeling endeavor involves preserving roto-translational invariance to ensure that the generated molecular conformations adhere to the laws of physics. Our proposed framework adeptly navigates this complexity, facilitating comprehensive training from inception to conclusion through the fine-tuning of a weighted variational lower bound that considers the conditional probability of conformations. Extensive experimental evaluations confirm the efficacy of the proposed DDGF in comparison to state-of-the-art methods.
@@ I Introduction
Graph representation learning has achieved remarkable success in modeling molecules for various tasks, encompassing property prediction [1, 2], graph domain adaptation [3, 4, 5], time series prediction [6, 7], and molecule generation [8, 9]. In this approach, molecules are typically represented as atom-bond graphs. Although this representation proves effective in multiple applications, a more intrinsic and informative depiction of molecules resides in their 3D geometry, often referred to as conformation. In this conformation, atoms are depicted using their Cartesian coordinates. The 3D structures exert significant influence over the biological and physical characteristics of molecules, rendering them crucial in various domains such as computational drug and material design  [10, 11, 12, 13]. Despite these advancements, predicting stable molecular conformations remains a formidable challenge. Conventional approaches reliant on molecular dynamics (MD) or Markov chain Monte Carlo (MCMC) methods are computationally intensive and struggle to efficiently handle large molecules [14]. To address this challenge, researchers are exploring novel methods that integrate graph-based models with physics-based simulations or leverage reinforcement learning techniques to optimize molecular conformations.
@@
Experimentally determining the three-dimensional arrangements of molecules can be a challenging and resource-intensive endeavor, characterized by high costs and lengthy timeframes. As a result, the pursuit of computationally deriving precise and energetically favorable molecular structures has gained prominence within the field of computational chemistry. Traditional approaches in this domain have traditionally relied on methods such as Markov chain Monte Carlo (MCMC) or molecular dynamics (MD) simulations [15]. These methods involve proposing potential conformations and assigning energy levels to them through either cost-effective empirical potential models or high-precision quantum chemical calculations  [16]. Recently, there has been a shift towards the integration of machine learning algorithms to estimate the conditional probability distribution $p(\mathbf{C}|G)$ for generating stable conformations $\mathbf{C}$ of a given molecular graph $G$. This approach builds upon datasets of molecules with known stable conformations  [17, 18]. Notably, two pioneering studies [18, 19] have introduced a sequential methodology. The initial phase involves predicting interatomic distances, which lays the foundation for the subsequent creation of molecular structures based on these predicted distances. This task is facilitated by solving a set of distance geometry challenges [20]. By leveraging the principles of distance geometry, these innovative methods successfully capture the essential characteristics of rotation and translation invariance in molecular structures, leading to increasingly accurate predictions of molecular conformations.
@@
The sequential methodologies that initially estimate interatomic distances and then deduce conformations are susceptible to a notable drawback: the potential for inaccuracies in local three-dimensional geometries. There are instances where the models predicting distances may inadvertently endorse geometrically incompatible sets of distances, leading to significant deviations. These inaccuracies, when fed into the conformation generation stage, can produce atypical and implausible molecular structures. The source of such discrepancies can often be attributed to the distance prediction models, which prioritize maximizing individual distance likelihoods rather than ensuring the holistic accuracy of the resulting conformations. To address this challenge, our research introduces a streamlined, end-to-end framework that directly formulates conformation predictions from the molecular graph, eliminating the intermediate distance prediction phase. Our approach aims to comprehensively capture the intricate dependencies and interactions among atoms, thus striving to enhance the precision and reliability of conformation prediction. This shift towards an integrated predictive model aligns with the evolution observed in the field of protein structure prediction. Innovations exemplified by the AlphaFold2 algorithm underscore the value and impact of adopting such an end-to-end approach. AlphaFold2’s improvements over its predecessor, AlphaFold, highlight the advantages of consolidating distinct phases into a cohesive modeling process  [21, 22]. Although AlphaFold2’s internal mechanisms are not fully disclosed, its remarkable success in deducing protein structures from linear sequences underscores the potential of end-to-end methodologies in complex structural prediction tasks.
@@
In our research, we introduce DDGF, an advanced and integrated framework specifically designed for the synthesis of molecular conformations using a sophisticated bilevel programming approach. A key aspect of our method is the incorporation of roto-translational invariance within the conformational modeling scheme, where interatomic distances play a pivotal role. Our innovative approach goes beyond traditional distance prediction models that aim to minimize prediction errors in the distance space. By framing the problem within the context of bilevel optimization [23, 24], DDGF simultaneously addresses both the prediction of distances and the resolution of distance geometry challenges that are fundamental to conformational synthesis. To the best of our knowledge, DDGF is the first algorithm in the field of molecular conformation synthesis that is trainable end-to-end while preserving the crucial aspects of roto-translational invariance. Our extensive experiments demonstrate the exceptional capabilities of DDGF, consistently outperforming the current leading methods across a range of well-established benchmarks covering conformational synthesis and distance distribution analysis. Furthermore, our findings underscore the importance of an end-to-end training objective in the generation of conformations that are not only realistic but also chemically meaningful.
@@ II Related Work
Conformation generation. Recent advancements in computational molecular modeling have given rise to a multitude of innovative deep generative models and methodologies, each presenting its unique advantages and inherent challenges. For instance, the CVGAE [17] harnessed a variational autoencoder (VAE) for the direct computation of 3D atomic structures. However, this model faced limitations in adequately preserving the crucial property of roto-translation equivariance, ultimately resulting in suboptimal performance. In response to this limitation, newer models have turned to intermediate structural descriptors, such as atomic distances and dihedral angles, known for their roto-translationally invariant nature, which is essential for accurately representing molecular shapes[25]. Yet, this modular approach, relying on these intermediate descriptors, introduces certain inefficiencies and complexities during both the training phase and sample generation. Approaches like those proposed by [18] and [19] involve predicting interatomic distance matrices and subsequently applying Distance Geometry solutions to derive spatial coordinates  [20]. Although CONFVAE represented a step forward with its unified, bilevel optimization-based end-to-end framework  [26], these techniques still grapple with the issue of error magnification. Inaccuracies in distance estimations often lead to misguidance of the coordinate determination mechanisms, resulting in the creation of molecular structures that are not only inaccurate but sometimes structurally implausible.
@@
In an attempt to tackle this challenge, the CONFGF model, as introduced in works by [27, 28], aimed to learn the gradient of log-likelihood concerning coordinates. Nevertheless, in practical applications, the model still depended on intermediate geometric elements. It estimated the gradient with respect to interatomic distances using denoising score matching (DSM) [29, 30] and subsequently applied the chain rule to derive the gradient of coordinates. Unfortunately, by utilizing DSM (Distance Gradient-based learning), the model was trained using perturbed distance matrices, which had the potential to violate the triangular inequality or contain negative values. Consequently, the model acquired knowledge from these incorrect distance matrices but was evaluated using valid ones computed from coordinates. This discrepancy between training and testing data resulted in a significant out-of-distribution issue  [31]. A recent approach called GEOMOL [32] introduced a highly systematic rule-based pipeline. It primarily focused on predicting a minimal set of geometric quantities, such as lengths and angles, and then reconstructed both local and global structures of conformations through a sophisticated procedure. Additionally, there have been efforts to leverage reinforcement learning for conformation search, exemplified by methods like TorsionNet [33]. However, it’s important to note that this approach fundamentally differs from other existing methodologies as it relies on the rigid rotor approximation and only models torsion angles.
@@
Advances in Protein Topology Prediction. The field of protein topology prediction has experienced rapid expansion, with a multitude of innovative methodologies emerging. These innovations encompass a range of approaches, including the use of flow-based models [34, 26], which effectively map the conformations of protein main chains. Additionally, techniques employing recurrent neural architectures have been employed for the sequential modeling of amino acids  [35].  [36] introduced a novel approach that utilizes neural networks to emulate an energy landscape, facilitating the inference of protein folds. The groundbreaking AlphaFold algorithm [21, 37] has significantly raised the bar in terms of accuracy for protein structure prediction. It is important to highlight, however, that proteins are structurally simple, primarily linear molecules, which is in contrast to the complex and branched configurations often found in general molecules, including various ring structures. This structural disparity makes the direct application of protein-folding methodologies less suitable for the intricate task of predicting the conformation of general molecular structures. Hence, there is a compelling need for distinct modeling techniques that are tailored to address the unique challenges posed by non-protein molecular structures.
@@
Point cloud generation. In the domain of three-dimensional structural synthesis, diffusion-driven models have emerged as significant contributors, as evidenced by pioneering efforts in recent literature [38, 39]. These approaches have primarily focused on the generation of point clouds, which represents a distinct area of study from our current research. Point clouds typically lack the graph-like structure that encompasses diverse atomic and bonding information, which is a fundamental aspect of our model’s domain. Moreover, the principle of equivariance, although not extensively incorporated in these point cloud methodologies, constitutes a cornerstone in our approach. This disparity in foundational concepts underscores the novelty of our model when compared to the landscape of existing methods.
@@
In conclusion, the realm of deep generative models has ushered in a new era for the generation of molecular conformations. Each innovation brings unique advantages while also encountering specific challenges. There remains an ongoing necessity for exploration and advancement to overcome the existing hurdles in improving the accuracy and computational efficiency of these conformational generation techniques.
@@ III-D Bilevel Optimization
Bilevel programs are a class of optimization problems in which a group of variables participating in the primary (outer) objective function are determined through the solution of a secondary (inner) optimization problem. To define this formally, we have the outer objective function denoted as F, the inner objective function as H, and the respective outer and inner variables represented as $\theta$ and $w$, a bilevel program can be formulated by:
@@
Bilevel programs have proven to be effective in various scenarios, including hyperparameter optimization, adversarial and multi-task learning, as well as meta-learning.
@@
Here, the symbol $\partial$ represents the partial derivative used for computing the Jacobian with respect to immediate variables, whereas $\nabla$ signifies a total derivative that considers the recursive calls made to the function F. The gradient mentioned above can be efficiently computed by unfolding the optimization dynamics using back-propagation, which is essentially reverse mode automatic differentiation [45]. We repeatedly substitute $w_{\Phi,t}=\Phi(w_{\theta,t-1},\theta)$ and apply the chain rule.
@@ IV-A Diffusion Process
Drawing inspiration from thermodynamics, we conceptualize the conformations of molecules, denoted as C, as particles within a constantly fluctuating system. Over time, the initial equilibrium state $\textbf{c}^{0}$ undergoes a stochastic journey, progressing through a series of increasingly disordered states $\textbf{c}^{t}$, until it reaches a state of maximal entropy resembling white noise after $T$ transitional steps. This unique progression, distinct from conventional latent variable constructs, is a defining feature of our diffusion model. It is characterized by a predetermined posterior distribution, denoted as $q(\mathbf{c}^{1:T}|\mathbf{c}^{0})$, which remains constant throughout the learning process. Specifically, we define this distribution as a Markov chain governed by a fixed variance schedule $\beta_{1},\cdots,\beta_{T}$:
@@
It’s important to note that in this work, we do not enforce specific invariance requirements on the diffusion process. The primary criterion is its efficiency in generating noisy samples for the training of the generative process, denoted as $p_{\theta}(\mathbf{c}^{0})$.
By defining $\alpha_{t}=1-\beta_{t}$ and ${\bar{\alpha}}_{t}=\Pi^{t}_{s=1}\alpha_{s}$, we can observe a significant aspect of the forward process: the conditional distribution $q(\mathbf{c}^{t}|\mathbf{c}^{0})$ at any time step $t$ can be precisely expressed as $q(\mathbf{c}^{t}|\mathbf{c}^{0})=\mathcal{N}(\mathbf{c}^{t};\sqrt{{\bar{\alpha}}_{t}}\mathbf{c}^{0},(1-{\bar{\alpha}}_{t})I)$. When we extend this process adequately up to $T$ steps, it inherently transforms the initial state $\mathbf{c}^{0}$ into a standard Gaussian distribution. This implies that a suitable choice for the prior distribution $p(\mathbf{c}^{T})$ would be an isotropic Gaussian distribution.